# Spatio-Temporal Adaptive Diffusion Models for EEG Super-Resolution in Epilepsy Diagnosis

### v3

@@ 2.3.3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like epileptic source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [44, 45, 46], the MTDM employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks [47] are introduced to capture multi-scale temporal features in the reverse denoising process .
@@
As shown in the blue part of Fig. 1, the MTDM mainly consists of position encoding, multi-scale 1D convolution blocks and diffusion Transformer blocks. Each diffusion Transformer block includes a normalization layer, a MSA layer, a cross-attention layer and a feed-forward network. The MTDM $\epsilon_{\theta}$ takes $z_{t}$, timestep $t$ and spatio-temporal condition $c$ as inputs, outputting the predicted noise $\epsilon_{\theta}(z_{t},t,c)$ at timestep $t$. The multi-scale 1D convolution blocks extract temporal features at different scales from the latent vectors $z_{t}$ using 1D convolution layers with various kernel sizes and concatenate them along the feature dimension. This convolution path yields an output that can be represented as:
@@
$\widetilde{H}_{t}\ =Concat(\widetilde{h}_{1},\ldots,\widetilde{h}_{n})$
@@
Initialize parameters $\theta$ of the STADMs
@@
      Perform: Extract spatio-temporal features of LR EEG $x$ via STCM ;
@@
      Perform: Reverse Denoising Process via MTDM $\epsilon_{\theta}$ ;
@@
Algorithm 1 Training Phase of STADMs
2.4 Loss Function
$+\mathrm{KL}(q(z_{t-1}|z_{0},z_{t})||p(z_{t-1}))$
where KL denotes the Kullback-Leibler divergence. However, directly predicting $z_{t-1}$ may lead to a complex and unstable training process. To simplify the training process, the MTDM $\epsilon_{\theta}$ is employed to predict the noise at each time step $t$, and the MSE loss is minimized to learn the model parameters between the predicted noise $\epsilon_{\theta}(z_{t},t,\tau_{\theta}(x))$ and actual noise $\epsilon$, where $x$ denotes the input LR EEG. The final loss function can be expressed as:
where $\epsilon_{\theta}$ is the proposed multi-scale Transformer denoising model and $\tau_{\theta}$ represents the proposed STCM. Algorithm 1 presents the entire training procedure along with the final loss function.
3 Experiment
3.1 Dataset and Preprocessing
In this work, we leveraged a publicly accessible dataset (Localize-MI [49]) for model training and validation. The dataset consists of high-density EEG data collected from seven drug-resistant epilepsy patients (mean age=35.1, $sd$ age=5.4), comprising a total of 61 sessions (average session duration per subject=8.71, $sd$ session duration per subject=2.65). Additionally, the dataset contains the spatial locations of stimulus contacts in MRI space, Freesurfer surface space [50], and MNI152 space [51]for each subject, along with the digital coordinates of 256 scalp EEG electrodes. We employed the MNE software for preprocessing of the raw data. Following the work of [49], we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) at a sampling rate of 8000 Hz for 256-channel EEG data, and deployed with the Geodesic Sensor Net using the HydroCel CleanLead electrode distribution system.
Table 1: The corresponding electrode systems for different scaling factors.
@@ 3.1 Dataset and Preprocessing
In the experiment, we downsampled the original HR EEG data along the channel dimension to obtain LR counterparts at different spatial resolutions, with scaling factors approximately set to 2, 4, 8, and 16. To simulate the acquisition of low-density EEG data in the real world, we mapped each scaling factor to different numbers of channels in the EGI standard electrode systems[52], as shown in Table 1. This process involved obtaining corresponding EEG electrode arrays. By using electrode systems with different channel ranks, we identified and removed the channels to be estimated, resulting in LR EEG that matched the preset channel levels. These LR-HR data pairs at different spatial resolutions were then created for model training.
3.2 Experiment Settings
3.2.1 Implement Details
The proposed STADMs are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. With the Adam optimizer, training of the model was carried out using a learning rate of $2\times 10^{-4}$, 300 epochs, and a batch size of 32. In the diffusion model framework, the forward process was set to 1000 time steps, with noise levels following sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with respective kernel sizes of 3, 5, 7, and 9. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
3.2.2 Metrics
To evaluate the quality of the synthetic EEG signals, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Normalized Mean Squared Error (NMSE), Signal-to-Noise Ratio (SNR) [53], and Mean Absolute Error (MAE). These metrics offer a thorough understanding of the accuracy and reliability of the synthetic signals in comparison to the real signals.
@@ 3.2.2 Metrics
3.3 Evaluation of EEG SR Reconstruction Performance
In this section, the performance of the proposed STADMs for EEG spatial super-resolution reconstruction is evaluated through various quantitative metrics and scalp topography maps.
3.3.1 Comparison With Other Existing EEG SR Methods
the proposed STADMs are compared with the existing EEG SR methods to demonstrate the effectiveness of the synthetic HR EEG (256 channels). We quantitatively analyze the reconstruction performance differences between the proposed STADMs and several other methods: CNN-based methods (Deep-CNN [18]), autoencoder-based methods [Weighted gate layer autoencoders (WGLAE) [28], Deep autoencoder (DAE) [54]], and GANs-based method (EEGSR-GAN) [29].
Fig. 2 shows quantitative comparison results between different methods. As indicated, the proposed STADMs exhibit superior performance, significantly surpassing other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STADMs maintain high signal quality and correlation. Moreover, STADMs significantly outperform other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STADMs and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG SR methods, the proposed STADMs effectively reconstruct high spatial resolution(256 channels) EEG, achieving significant spatial resolution enhancement of LR EEG.
@@ 3.3.1 Comparison With Other Existing EEG SR Methods
Figure 2: Quantitative comparison results of STADMs and existing EEG SR methods using four metrics. (a): The PCC performance across various methods. (b): The MAE performance across various methods. (c): The NMSE performance across various methods. (d): The SNR performance across various methods. The synthetic results of STADMs demonstrate the highest quality across these four quantitative evaluation metrics.
3.3.2 Comparison With Different scaling factors
To evaluate the performance of the proposed STADMs in generating synthetic SR EEG from LR EEG under different scaling factors, we employ four metrics for quantitative analysis of the model’s generative performance. Fig. 3 shows quantitative results for different scaling factors. As the scaling factor increases, PCC and SNR values decrease, indicating a weakening correlation between synthetic SR EEG and ground truth. Meanwhile, MAE and NMSE values increase, implying a growing difference. This suggests that the reconstruction performance of STADMs diminishes at larger scaling factors due to limited channel information, making HR EEG reconstruction challenging. However, even at larger scaling factors, the performance of STADMs remains comparable to that at smaller scaling factors. For example, with a scaling factor of 16, the PCC value is only 3.02% lower than at a scaling factor of 8, the MAE value differs by 1.68%, and the SNR value by 3.19%. This indicates that STADMs can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STADMs effectively enhance LR EEG spatial resolution across different scaling factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
@@ 3.3.2 Comparison With Different scaling factors
3.3.3 Qualitative results of Synthetic EEG
@@
3.4 Classification Evaluation on the Synthetic SR EEG
@@
Table 2: The Clissfication Performance Comparison of Different Frequency Bands.
@@
Fig. 5 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STADMs outperforms LR EEG in all metrics, showing that STADMs enhance classification performance of EEG-Net. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STADMs effectively enhance LR EEG spatial resolution, improving downstream classification tasks. Moreover, this provides a novel approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to $c\times 5$ features (where c denotes the number of channels), and classification is performed based on these features. Table 2 shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands ($\delta$, $\theta$, $\alpha$) have worse performance than high-frequency bands ($\beta$, $\gamma$), but the average accuracy improvement in low-frequency bands (5.06%) surpasses that in high-frequency bands (4.95%). This suggests that the proposed STADMs more Adequately preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
3.5 Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [57]. To determine the impact of the proposed STADMs in enhancing EEG spatial resolution and its effectiveness in improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyses. We utilize localization error [7] for quantitative analysis, defining it as the Euclidean distance between the source localization results and the ground-truth positions. This metric is employed to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [58], an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [59]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
@@ 3.5 Source Localization Evaluation on the Synthetic EEG
Fig. 6 shows a comparison of localization errors for different EEG types among all subjects in the dataset. The localization error for LR EEG is widely dispersed with higher values, indicating lower precision and greater variability. In contrast, SR EEG shows significantly lower and more concentrated localization errors, suggesting higher precision and stability. Although the localization error of SR EEG is slightly higher than HR EEG, its accuracy is markedly improved, making it a viable substitute for HR EEG and reducing reliance on expensive high-density equipment. Therefore, the experimental results demonstrate that synthetic SR EEG substantially enhances the spatial resolution and source localization accuracy of LR EEG.
@@
Fig. 7 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results, and blue circles show the results from the LR EEG. The LR EEG results display considerable deviations from the ground truth, while the SR EEG results show minimal deviation, closely aligning with HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, the method not only enhances EEG spatial resolution but also supports high-precision EEG imaging analysis in clinical and research applications.
4 Discussion and Conclusion
@@
In this work, STADMs are proposed to address the spatial resolution limitations of low-density EEG devices, particularly in clinical diagnostic applications, such as epilepsy focus localization. STADMs are the first to employ a diffusion model to achieve spatial SR reconstruction from LR EEG to HR EEG. To ensure that the generated results align with subject-specific characteristics, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which are then used as conditional inputs to guide the reverse denoising process. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance, bridging the gap between SR EEG and HR EEG. Qualitative and quantitative results indicate that STADMs effectively enhance the spatial resolution of EEG. Furthermore, classification and source localization experiments demonstrate their ability to improve EEG performance in practical scenarios.
@@
Experimental results show that while the proposed STADMs outperform existing methods in reconstruction performance, larger scaling factors reveal a significant difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 3. We identified several possible reasons for this: i) With larger scaling factors, the available spatial information for the model is constrained, making it challenging to accurately capture the spatial relationships between channels. ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model. iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of other physiological signals. In the future, we will further optimize this framework and apply it to practical clinical scenarios in epilepsy diagnosis, aiding doctors in the preoperative assessment for epilepsy focus resection surgery.

### v4

@@ III-C3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [39], the MTD employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks are introduced to capture multi-scale temporal features in the reverse denoising process.
@@
As shown in the blue part of Fig. 2, the MTD mainly consists of position encoding, multi-scale 1D convolution blocks and diffusion Transformer blocks. Each diffusion Transformer block includes a normalization layer, a MSA layer, a cross-attention layer and a feed-forward network. The MTD $\epsilon_{\theta}$ takes $z_{t}$, timestep $t$ and spatio-temporal condition $c$ as inputs, outputting the predicted noise $\epsilon_{\theta}(z_{t},t,c)$ at timestep $t$. The multi-scale 1D convolution blocks extract temporal features at different scales from the latent vectors $z_{t}$ using 1D convolution layers with various kernel sizes and concatenate them along the feature dimension. This convolution path yields an output that can be represented as:
@@
$\tilde{H}_{t}\ =Concat(\widetilde{h}_{1},\ldots,\widetilde{h}_{n})$
@@
III-D Loss Function
$+\mathrm{KL}(q(z_{t-1}|z_{0},z_{t})||p(z_{t-1}))$
where KL denotes the Kullback-Leibler divergence. However, directly predicting $z_{t-1}$ may lead to a complex and unstable training process. To simplify the training process, the MTD $\epsilon_{\theta}$ is employed to predict the noise at each time step $t$, and the MSE loss is minimized to learn the model parameters between the predicted noise $\epsilon_{\theta}(z_{t},t,\tau_{\theta}(x))$ and actual noise $\epsilon$, where $x$ denotes the input LR EEG. The final loss function can be expressed as:
where $\epsilon_{\theta}$ is the proposed multi-scale Transformer denoising model and $\tau_{\theta}$ represents the proposed STC. Algorithm 1 presents the entire training procedure along with the final loss function.
@@ III-D Loss Function
Initialize parameters $\theta$ of the STAD
@@
      Perform: Extract spatio-temporal features of LR EEG $x$ via STC ;
@@
      Perform: Reverse Denoising Process via MTD $\epsilon_{\theta}$ ;
@@
Algorithm 1 Training Phase of STAD
IV Experiment
IV-A Dataset and Preprocessing
In this work, the EEG dataset is collected from a publicly accessible dataset (Localize-MI [45]), which consists of high-density EEG data collected from seven drug-resistant epilepsy patients (mean age=35.1, $sd$ age=5.4), comprising a total of 61 sessions (average session duration per subject=8.71, $sd$ session duration per subject=2.65). Additionally, the dataset contains the spatial locations of stimulus contacts in MRI space, Freesurfer surface space [46], and MNI152 space [47]for each subject, along with the digital coordinates of 256 scalp EEG electrodes.
TABLE I: The corresponding electrode systems for different scaling factors.
@@ IV-A Dataset and Preprocessing
We employed the MNE software for preprocessing of the raw data. In this work, we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) [48] at a sampling rate of 8000 Hz for 256-channel EEG data. Besides, we downsampled the original HR EEG data along the channel dimension to obtain LR counterparts at different spatial resolutions, as shown in Table I. By using electrode systems with different channel ranks, we identified and removed the channels to be estimated, resulting in LR EEG that matched the preset channel levels.
IV-B Experiment Settings
IV-B1 Implement Details
The proposed STAD are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. With the Adam optimizer, training of the model was carried out using a learning rate of $2\times 10^{-4}$, 300 epochs, and a batch size of 32. In the diffusion model framework, the forward process was set to 1000 time steps, with noise levels following sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with respective kernel sizes of 3, 5, 7, and 9. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
IV-B2 Metrics
To evaluate the quality of the synthetic EEG signals, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Normalized Mean Squared Error (NMSE), Signal-to-Noise Ratio (SNR), and Mean Absolute Error (MAE). These metrics offer a thorough understanding of the accuracy and reliability of the synthetic signals in comparison to the real signals.
@@ IV-B2 Metrics
IV-C Evaluation of EEG SR Reconstruction Performance
In this section, the performance of the proposed STAD for EEG spatial super-resolution reconstruction is evaluated through various quantitative metrics and scalp topography maps.
IV-C1 Comparison With Other Existing EEG SR Methods
the proposed STAD are compared with the existing EEG SR methods to demonstrate the effectiveness of the synthetic HR EEG (256 channels). We quantitatively analyze the reconstruction performance differences between the proposed STAD and several other methods: CNN-based methods (Deep-CNN [33]), autoencoder-based methods [Weighted gate layer autoencoders (WGLAE) [35], Deep autoencoder (DAE) [34]], and GANs-based method (EEGSR-GAN) [36].
@@ IV-C1 Comparison With Other Existing EEG SR Methods
Figure 3: Quantitative comparison results of STAD and existing EEG SR methods using four metrics. (a): The PCC performance across various methods. (b): The MAE performance across various methods. (c): The NMSE performance across various methods. (d): The SNR performance across various methods. The synthetic results of STAD demonstrate the highest quality across these four quantitative evaluation metrics.
Fig. 3 shows quantitative comparison results between different methods. As indicated, the proposed STAD exhibit superior performance, significantly surpassing other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STAD maintain high signal quality and correlation. Moreover, STAD significantly outperform other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STAD and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG SR methods, the proposed STAD effectively reconstruct high spatial resolution(256 channels) EEG, achieving significant spatial resolution enhancement of LR EEG.
IV-C2 Comparison With Different scaling factors
@@
To evaluate the performance of the proposed STAD in generating synthetic SR EEG from LR EEG under different scaling factors, we employ four metrics for quantitative analysis of the model’s generative performance. Fig. 4 shows quantitative results for different scaling factors. As the scaling factor increases, PCC and SNR values decrease, indicating a weakening correlation between synthetic SR EEG and ground truth. Meanwhile, MAE and NMSE values increase, implying a growing difference. This suggests that the reconstruction performance of STAD diminishes at larger scaling factors due to limited channel information, making HR EEG reconstruction challenging. However, even at larger scaling factors, the performance of STAD remains comparable to that at smaller scaling factors. For example, with a scaling factor of 16, the PCC value is only 3.02% lower than at a scaling factor of 8, the MAE value differs by 1.68%, and the SNR value by 3.19%. This indicates that STAD can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STAD effectively enhance LR EEG spatial resolution across different scaling factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
IV-C3 Qualitative results of Synthetic EEG
@@
IV-D Classification Evaluation on the Synthetic SR EEG
@@
Fig. 6 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STAD outperforms LR EEG in all metrics, showing that STAD enhance classification performance of EEG-Net. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STAD effectively enhance LR EEG spatial resolution, improving downstream classification tasks. Moreover, this provides a novel approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to $c\times 5$ features (where c denotes the number of channels), and classification is performed based on these features. Table II shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands ($\delta$, $\theta$, $\alpha$) have worse performance than high-frequency bands ($\beta$, $\gamma$), but the average accuracy improvement in low-frequency bands (5.06%) surpasses that in high-frequency bands (4.95%). This suggests that the proposed STAD more Adequately preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
TABLE II: The Clissfication Performance Comparison of Different Frequency Bands.
@@
IV-E Source Localization Evaluation on the Synthetic EEG
@@
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [51]. To determine the impact of the proposed STAD in enhancing EEG spatial resolution and its effectiveness in improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyses. We utilize localization error [8] for quantitative analysis, defining it as the Euclidean distance between the source localization results and the ground-truth positions. This metric is employed to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [52], an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [53]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
Fig. 7 shows a comparison of localization errors for different EEG types among all subjects in the dataset. The localization error for LR EEG is widely dispersed with higher values, indicating lower precision and greater variability. In contrast, SR EEG shows significantly lower and more concentrated localization errors, suggesting higher precision and stability. Although the localization error of SR EEG is slightly higher than HR EEG, its accuracy is markedly improved, making it a viable substitute for HR EEG and reducing reliance on expensive high-density equipment. Therefore, the experimental results provide compelling evidence that the application of super-resolution techniques to LR EEG data substantially enhances the spatial resolution and source localization accuracy.
@@
Fig. 8 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results, and blue circles show the results from the LR EEG. The LR EEG results display considerable deviations from the ground truth, while the SR EEG results show minimal deviation, closely aligning with HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, By leveraging this method to augment the spatial resolution of LR EEG, researchers and clinicians can overcome the limitations imposed by low-density electrode configurations, facilitating more precise localization of neural sources and enabling a deeper understanding of brain dynamics.
V Discussion and Conclusion
@@
In this work, STAD are proposed to address the spatial resolution limitations of low-density EEG devices, particularly in clinical diagnostic applications, such as epilepsy focus localization. STAD are the first to employ a diffusion model to achieve spatial SR reconstruction from LR EEG to HR EEG. To ensure that the generated results align with subject-specific characteristics, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which are then used as conditional inputs to guide the reverse denoising process. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance, bridging the gap between SR EEG and HR EEG. Qualitative and quantitative results indicate that STAD effectively enhance the spatial resolution of EEG. Furthermore, classification and source localization experiments demonstrate their ability to improve EEG performance in practical scenarios. In summary, By bridging the gap between low-cost, low-density EEG systems and the spatial resolution requirements of emerging neurotechnological applications, SR EEG paves the way for more accessible and precise neuroimaging, ultimately advancing our understanding of brain function and facilitating the development of innovative neurological interventions.
@@
Experimental results show that while the proposed STAD outperform existing methods in reconstruction performance, larger scaling factors reveal a difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 4. We identified several possible reasons for this: i) With larger scaling factors, the available spatial information for the model is constrained, making it challenging to accurately capture the spatial relationships between channels. ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model. iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of EEG signals. In the future, we will further expand the application range of this framework to enhance other brain function signals. Through improving algorithms and models, we plan to apply this technology to more precise clinical diagnosis and treatment scenarios, such as BCI technology and neurofeedback, thus promoting research in neuroscience, cognitive science and other related fields and enhancing the understanding and diagnostic capabilities of neuropsychiatric diseases.